A 4.40 percent coupon municipal bond has 10 years left to maturity and has a price quote of 97.95. The bond can be called in four years. The call premium is one year of coupon payments. (Assume interest payments are semiannual and a par value of $5,000.)

Compute the bond’s current yield. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Compute the yield to maturity. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Compute the taxable equivalent yield (for an investor in the 35 percent marginal tax bracket). (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Compute the yield to call. (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

a)
Current yield = (Annual coupon amount / current price) *100
= (5,000 * 4.40%) / (5,000 * 97.95%) * 100
= (220 / 4,897.5) * 100
= 4.49%

b)
Yield to maturity can be calculated by using the following excel formula:
=RATE(nper,pmt,pv,fv)
=RATE(10,220,-4897.5,5000)
= 4.66%

c)
Taxable equivalent yield = Tax exempt yield / (1 - tax rate)
= 4.66 / (1 - 0.35)
= 7.17%

d)
Call price = 5,000 + 220 = 5220
Yield to call can be calculated by using the following excel formula:
=RATE(nper,pmt,pv,fv)
=RATE(4,220,-4897.5,5220)
= 6.00%